# Cancellation properties for exotic $4$-dimensional positive scalar curvature metrics

### v2

@@ 4.3. Construction of the invariants and their evaluation
We now proceed exactly as in the case of $D$ to get the spectrum map $\sigma$. ∎
@@
There exist orientation-preserving diffeomorphisms $f_{j}:M_{j}\to M_{j}$ of $1$-connected $4$-manifolds, $j=1,2,3$, such that $\{\varphi(\tilde{\alpha}(f_{1})),\varphi(\tilde{\alpha}(f_{2})),\varphi(\tilde{\alpha}(f_{3}))\}$ generates $\mathbb{Z}/2^{3}$. The manifolds $M_{1}$ and $M_{2}$ are spin.
@@
Let $f_{2}\in\mathrm{Diff}(S^{2}\times S^{2})$ be the product of two degree $-1$ diffeomorphisms. With respect to the above basis, we have
@@ 5. Proof of Theorem C
This is conceptually very similar, so we shall be quite sketchy. Let
@@
$\alpha_{*}:\pi_{j}(\mathrm{Diff}^{+}(M))=\pi_{j+1}(B\mathrm{Diff}^{+}(M))\to\pi_{j+1}(\Omega^{\infty}\mathrm{MT}\theta(4))\otimes\mathbb{Q}$
annihilates $x$, where again $\theta=\mathrm{SO}$ or $\mathrm{Spin}$, depending on the case at hand.
However, rational homotopy groups of $\Omega^{\infty}\mathrm{MT}\theta(4)$ are detected by rational Miller–Morita–Mumford classes (for the Euler class and the first Pontrjagin class), so it suffices to prove that the $M$-bundle $E_{x}\to S^{j+1}$ classified by $x$ does have trivial such classes. However, by construction, $E_{x}$ is trivial as a topological bundle. Since Miller–Morita–Mumford classes can be defined for bundles of topological manifolds by [9] (this is straightforward from Novikov’s theorem on the topological invariance of rational Pontrjagin classes), the claim follows.

### v3

@@ 4.3. Construction of the invariants and their evaluation
Using that $KO^{0}(B)=[B;\Omega^{\infty}\mathrm{KO}]$, we can view $\mathrm{ind}(D)$ and $\mathrm{ind}(D^{+})$ as a homotopy class of maps $B\to\Omega^{\infty}\mathrm{KO}$. Expressing these homotopy classes in topological terms is the aim of the Atiyah–Singer family index theorem [2], [3]. In [13], it was observed that the family index theorem can be reformulated in terms of Madsen–Tillmann spectra.
The operators $D$ and $D^{+}$ are “universal” in the sense of [13, §3.2], there, we constructed universal symbol classes $\mathrm{smb}_{D},\mathrm{smb}_{D^{+}}\in KO^{0}(\mathrm{MT}\mathrm{SO}(4))$ (denoted $\mathrm{th}\sigma_{D}$ in loc. cit.).
The classes $\mathrm{smb}_{D}$, $\mathrm{smb}_{D^{+}}$ are elements of the $K$-theory of the spectrum $\mathrm{MT}\mathrm{SO}(4)$ and can be viewed as maps $\mathrm{MT}\mathrm{SO}(4)\to\mathrm{KO}$ of spectra. Applying $\Omega^{\infty}$ produces maps $\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4)\to\Omega^{\infty}\mathrm{KO}$. As shown in [13, §3.2], the Atiyah–Singer family index theorem says that $(\Omega^{\infty}\mathrm{smb}_{D})\circ\alpha_{E}\sim\mathrm{ind}(D):B\to\Omega^{\infty}\mathrm{KO}$; and the analogous formula holds for $D^{+}$. We now put $\chi:=\mathrm{smb}_{D}$ and $\sigma:=\mathrm{smb}_{D^{+}}$.
The above outline was slightly inaccurate because [13] only covered the case of complex $K$-theory whereas here we need to use indices in real $K$-theory. One can modify the argument of that paper in order to cover the real case instead, using the real family index theorem [3] in place of [2] and Real $K$-theory [1]. Alternatively, the proof of the index theorem given in the more recent paper [11] deals with the real and complex case on the same footing. ∎
@@
There exist orientation-preserving diffeomorphisms $f_{j}:M_{j}\to M_{j}$ of $1$-connected 44-manifolds, $j=1,2,3$, such that $\{\varphi(\tilde{\alpha}(f_{1})),\varphi(\tilde{\alpha}(f_{2})),\varphi(\tilde{\alpha}(f_{3}))\}$ generates $\mathbb{Z}/2^{3}$. The manifolds $M_{1}$ and $M_{2}$ are spin.
@@
Let $f_{2}\in\mathrm{Diff}(S^{2}\times S^{2})$ be the product of two degree 1-1 diffeomorphisms. With respect to the above basis, we have
@@ 5. Proof of Theorem C
This is conceptually very similar to Theorem B. Let
@@
$\alpha_{*}:\pi_{j}(\mathrm{Diff}^{+}(M))=\pi_{j+1}(B\mathrm{Diff}^{+}(M))\to\pi_{j+1}(\Omega^{\infty}\mathrm{MT}\theta(4);\alpha(M))\otimes\mathbb{Q}$ (5.1)
annihilates $x$, where again $\theta=\mathrm{SO}$ or $\mathrm{Spin}$, depending on the case at hand. Let us first list recall the computation of rational homotopy and cohomology of $\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4)$ and $\Omega^{\infty}\mathrm{MT}\mathrm{Spin}(4)$; see [13, §2.4] for an exposition of this well-known material. We shall need the following facts.
  1. (1)
    The Thom isomorphism provides an isomorphism
    $\mathrm{thom}:H^{*}(B\mathrm{SO}(4);\mathbb{Q})\cong H^{*-4}(\mathrm{MT}\mathrm{SO}(4);\mathbb{Q})$
    of the cohomology of $B\mathrm{SO}(4)$ with the cohomology of the spectrum $\mathrm{MT}\mathrm{SO}(4)$ (this includes the case <4*<4). The cohomology suspension gives a map
    $\mathrm{susp}:H^{*}(\mathrm{MT}\mathrm{SO}(4);\mathbb{Q})\to H^{*}(\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4);\mathbb{Q})$
    to the cohomology of its infinite loop space. The cohomology of the infinite loop space is a graded commutative algebra, but the spectrum cohomology is only a graded vector space. The induced map of graded algebras
    $\Lambda(\mathrm{susp}\circ\mathrm{thom}):\Lambda(H^{*>4}(B\mathrm{SO}(4);\mathbb{Q})\to H^{*}(\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4);\mathbb{Q})$
    from the free graded commutative algebra on the source is an isomorphism.
  2. (2)
    For any class $c\in H^{*}(B\mathrm{SO}(4);\mathbb{Q})=\mathbb{Q}[e,p_{1}]$ of degree $k>4$, we obtain a class
    $\lambda_{c}:=\mathrm{susp}(\mathrm{thom}(c))\in H^{k-4}(\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4);\mathbb{Q}).$
    By step (1), $H^{*}(\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4);\mathbb{Q})$ is a polynomial algebra on all $\lambda_{c}$, where $c$ runs through all monomials on $e$ and $p_{1}$ of total degree >4>4.
  3. (3)
    If $\pi:E\to B$ is a bundle of closed oriented $4$-manifolds with Madsen–Tillmann map $\alpha_{E}:B\to\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4)$, we have
    $\alpha_{E}^{*}\lambda_{c}=\kappa_{c}(E):=\pi_{!}(c(T_{v}E)),$
    the Miller-Morita–Mumford class of $E$ associated to $c$, obtained by fibre integration of the characteristic class $c$ evaluated on the vertical tangent bundle.
  4. (4)
    Being a connected loop space, $\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4)$ has the property that the rational Hurewicz map $\pi_{*}(\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4))\otimes\mathbb{Q}\to H_{*}(\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4);\mathbb{Q})$ is injective. Hence a homotopy class $y\in\pi_{j}(\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4))$ vanishes after rationalization provided that $\langle\mathrm{hur}(y),x\rangle=0$ for all $x\in H^{*}(\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4);\mathbb{Q})$.
  5. (5)
    The same recipe applies to $\mathrm{MT}\mathrm{Spin}(4)$. Since the map $B\mathrm{Spin}(4)\to B\mathrm{SO}(4)$ induces an isomorphism in rational cohomology, the induced map
    $\Omega^{\infty}_{0}\mathrm{MT}\mathrm{Spin}(4)\to\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4)$
    induces an isomorphism in rational cohomology as well, and also on rational homotopy groups, by Serre class theory.
Proof of Theorem C.
Item (4) of the above leaves us to prove that for a $1$-connected 44-manifold $M$ and
$x\in\ker(\pi_{j}(\mathrm{Diff}^{+}(M))\to\pi_{j}(\mathrm{Homeo}^{+}(M)))$ (5.2)
($j\geq 0$), we have
$x\in\ker(\pi_{j}(\mathrm{Diff}^{+}(M))=\pi_{j+1}(B\mathrm{Diff}^{+}(M))\stackrel{{\scriptstyle\alpha_{*}}}{{\to}}\pi_{j+1}(\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4),\alpha(M))\otimes\mathbb{Q}).$ (5.3)
The image of $x$ under (5.3) is the Madsen–Tillmann map $\alpha_{E_{x}}:S^{j+1}\to\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4)$ of the smooth $M$-bundle $\pi:E_{x}\to S^{j+1}$ classified by $x$.
The target group of (5.3) is not taken at the basepoint; when composed with the isomorphism $\pi_{j+1}(\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4),\alpha(M))\otimes\mathbb{Q}\cong\pi_{j+1}(\Omega^{\infty}_{0}\mathrm{MT}\mathrm{SO}(4))\otimes\mathbb{Q}$ induced by the loop space structure, the map (5.3) takes $x$ to the difference of the Madsen–Tillmann map of $E_{x}$ and the Madsen–Tillmann map of the trivial bundle $p:S^{j+1}\times M\to S^{j+1}$ (compare the arguments leading to (4.17) above). Items (1)–(3) show that we must prove
$\kappa_{c}(E_{x})-\kappa_{c}(S^{j+1}\times M\to S^{j+1})\in H^{*}(S^{j+1};\mathbb{Q})$
for each $c\in H^{*}(B\mathrm{SO}(4);\mathbb{Q})$ of degree >4>4. Since $\kappa_{c}$ lives in positive degrees, the second summand vanishes, so we are left with $E_{x}$. By assumption (5.2), $E_{x}\to S^{j+1}$ is trivial as a bundle of topological manifolds. Since Miller–Morita–Mumford classes can be defined for bundles of topological manifolds by [12] (this is straightforward from Novikov’s theorem on the topological invariance of rational Pontrjagin classes), we obtain that $\kappa_{c}(E_{x})=0$, for all $c\in H^{*>4}(B\mathrm{SO}(4);\mathbb{Q})$. ∎